# Fractional Chern Insulator and Quantum Anomalous Hall Crystal in Twisted MoTe2

### v2

@@ Acknowledgements.
J.L.C., Q.Y.L., and W.L. are indebted to Bin-Bin Chen and Hongyu Lu for stimulating discussions. The iDMRG calculation is based on TeNPy [61]. The dynamical and finite-temperature calculations are based on TensorKit.jl [87] and FiniteMPS.jl [88]. This work was supported by the National Natural Science Foundation of China (Grant Nos. 12222412, 12047503), National Key Projects for Research and Development of China with Grant No. 2024YFA1409200, Innovation Program for Quantum Science and Technology (Nos. 2021ZD0301900), CAS Project for Young Scientists in Basic Research (Grant No. YSBR-057), and the Postdoctoral Fellowship Program of CPSF (Grant No. GZB20240772). We thank the HPC-ITP for the technical support and generous allocation of CPU time. X.Y.W. acknowledges financial support from the National High Magnetic Field Laboratory through National Science Foundation (NSF) Grant No. DMR-2128556, and the State of Florida.
@@ I.1 Continuum model of twisted bilayer MoTe2
$\begin{split}{H}_{\uparrow}=\int\mathrm{d}^{2}\mathbf{r}\begin{pmatrix}\psi^{\dagger}_{b\uparrow}(\mathbf{r})&\psi^{\dagger}_{t\uparrow}(\mathbf{r})\end{pmatrix}\begin{pmatrix}{h}_{b}({\mathbf{p}})+\Delta_{b}(\mathbf{r})+\frac{u_{D}}{2}&\Delta_{T}(\mathbf{r})\\ \Delta_{T}^{*}(\mathbf{r})&{h}_{t}({\mathbf{p}})+\Delta_{t}(\mathbf{r})-\frac{u_{D}}{2}\end{pmatrix}\begin{pmatrix}\psi_{b\uparrow}(\mathbf{r})\\ \psi_{t\uparrow}(\mathbf{r})\end{pmatrix}.\end{split}$ (S1)
Here $\psi^{\dagger}_{l\uparrow}(\mathbf{r})$ creates an electron in layer $l=b,t$ , with spin $\uparrow$ and in valley $\mathbf{K}$, and at position $\mathbf{r}$. For convenience we drop the valley subscript and label the electrons by their spin quantum number. ${h}_{l}({\mathbf{p}})=-({\mathbf{p}}-\hbar\mathbf{K}_{l})^{2}/2m^{*}$ is kinetic energy, with ${\mathbf{p}}=-i\hbar\nabla_{\mathbf{r}}$ and effective mass $m^{*}$. $\mathbf{K}_{l}$ points sit at the corners of the hexagonal moiré Brillouin zone (see Fig. S1(a) inset). We describe the effect of displacement field $\mathbf{D}$ by adding a potential difference $u_{D}$ between two layers. $\Delta_{b(t)}(\mathbf{r})=2V\sum_{j=1,3,5}\cos(\mathbf{G}_{j}\cdot\mathbf{r}\pm\phi)$ is the intralayer potential for bottom (top) layer, and $\Delta_{T}(\mathbf{r})=w(1+e^{-i\mathbf{G}_{2}\cdot\mathbf{r}}+e^{-i\mathbf{G}_{3}\cdot\mathbf{r}})$ is interlayer tunneling potential. $\mathbf{G}_{i=1,\dots 6}$ are the six smallest reciprocal lattice vectors, with $\mathbf{G}_{1}=\frac{4\pi}{\sqrt{3}L_{M}}\hat{x}$ along the $+x$ direction, and $\mathbf{G}_{i+1}$ is the vector obtained by rotating $\mathbf{G}_{i}$ $60^{\circ}$ counterclockwise. Here $L_{M}\approx a_{0}/\theta$ is the moiré unit vector length, with $\theta$ the twist angle and $a_{0}$ the MoTe2 lattice spacing.
$a_{0}$ [Å] $m^{*}$ [$m_{e}$] $V$ [meV] $\phi$ $\omega$ [meV]
Ref. [6] 3.47 0.62 8.0 -$89.6^{\circ}$8.5
Ref. [53] 3.55 0.62 17.0 $107.7^{\circ}$16.0
Ref. [18] 3.52 0.6 20.8 $107.7^{\circ}$23.8
Ref. [19] 3.52 0.62 7.5 -$100^{\circ}$11.3
@@ I.2 Tight binding fit to the continuum model
With the exception of Ref. [19], other DFT calculations predict the upper two moiré bands of a given valley carry Chern numbers $+1$ and $-1$ respectively, including the one we used for this work. This allows us to construct exponentially localized Wannier orbitals without suffering from topological obstruction of a band composite with a finite Chern number [57]. Using the trial state procedure discussed in Ref. [57], we construct exponentially localized Wannier orbitals to describe the Hilbert space spanned by the upper two valence bands per valley, see Fig. S1(b). The two Wannier orbitals per valley reside predominantly on the honeycomb lattice, which is the dual lattice of the moiré triangular lattice made of AA stackings. We hereby label these states as $w_{A}(\mathbf{r})$ and $w_{B}(\mathbf{r})$ respectively, with the subscripts $A,B$ denoting the center of the Wannier orbitals. Separately, these two orbitals are layer polarized, with $w_{A(B)}(\mathbf{r})$ having most orbital weights on the bottom (top) layer.
@@ I.3 Screened Coulomb interaction
We note a key distinction here for tMoTe2 as compared to the twisted bilayer graphene (TBG): due to the nontrivial topology in TBG, it is not possible to construct exponentially localized and symmetric Wannier orbitals [57, 90, 91, 92, 93], and the Coulomb interaction projected onto the narrow bands can have additional contributions beyond the density-density interaction in Eq. (S10), notedly the assisted hopping terms [94, 95, 71]. However, for tMoTe2 the Wannier orbitals are sufficiently localized on the honeycomb lattice sites (see Fig. S1b), and as a result, the interaction is effectively of the Coulomb density-density form [23].
@@ I.4 Particle-hole transformation
Alternatively, we can perform a particle-hole transformation $d^{\dagger}_{i\sigma}=c_{i\sigma}$ which inverts the non-interacting band structure (as well as the hopping parameters and chemical potential), while simultaneously we preserve the same functional form for the projected density density interactions but with $n_{i\sigma}=d^{\dagger}_{i\sigma}d_{i\sigma}$. Such a transformation has been used in previous theoretical studies on tMoTe2 (see, e.g., Ref. [19]). Here $|{\Omega}\rangle$ will be a vacuum state for $d$ fermions, and therefore the density-density interaction will not lead to band renormalizations.
Therefore, instead of stating our results for hole fillings $-1\leq\nu_{h}<0$, we report results for electron fillings $0<\nu\leq 1$. Accordingly, we present the band structure of the model described by Eq. (S2) in Fig. S3, which includes hopping terms up to the third nearest neighbors following the particle-hole transformation. Without interactions, electrons only fill the lower band [see, Fig. S3(c)] when filling factor $\nu\leq 1$.
II Supplementary ground-state results
II.1 Magnetic gaps for various fillings
@@
II.2 Ground-state results for $\nu=1$
@@
II.3 Ground-state results for fillings $\nu=p/q$ with $q\geq 5$
@@
II.3.1 Charge pumping results
@@ II.3.3 Phase diagram
III Supplementary finite-temperature results
Figure S9: Finite-temperature results for an XC4×12×241224$\times$ 12$\times$ 24 × 12 × 2 with parameters $\nu=1$ and $\epsilon_{r}=15$. (a) Electron density distribution $n_{\uparrow(\downarrow)}(x)$ at $T\approx 1.45\textrm{ K}$. (b) Temperature evolution of the bulk-averaged double occupancy $D_{n}$. The inset shows $\partial D_{n}/\partial\ln T$, which determines a characteristic temperature $T_{d}^{\ast}\approx 662$ K where $D_{n}$ is suppressed rapidly. (c,d) Comparison of low-temperature electron specific heat $c_{e}$ and magnetic moment $m^{z}$ per electron for different pinning fields $h^{z}$. Panels (a,b) and (c,d) are computed with bond dimension $D=2000$ and $D=3000$, respectively.
@@ III Supplementary finite-temperature results
IV computation method and convergence check of the spectral function
@@
It remains to compute the expectation values like $GS\expectationvalue{c_{\mathbb{k}\alpha}e^{-iHt}c_{\mathbb{k}\beta}^{\dagger}}{\rm GS}$ and $GS\expectationvalue{c_{\mathbb{k}\alpha}^{\dagger}e^{-iHt}c_{\mathbb{k}\beta}}{\rm GS}$. Considering the former one, it can be written as the inner product of two states $e^{iHt/2}c_{\mathbb{k}\alpha}^{\dagger}|{\rm GS}\rangle$ and $e^{-iHt/2}c_{\mathbb{k}\beta}^{\dagger}|{\rm GS}\rangle$. We first use standard DMRG algorithm to obtain the ground state $|{\rm GS}\rangle$ that is represented as a MPS. Then, we act the operator $c_{\mathbb{k}\alpha}^{\dagger}$, which can be efficiently represented as a MPO with bond dimension $D=2$ via automata, on the ground state MPS and obtain the new MPS $c_{\mathbb{k}\alpha}^{\dagger}|{\rm GS}\rangle$. $c_{\mathbb{k}\beta}^{\dagger}|{\rm GS}\rangle$ is obtained similarly. Then, we use the TDVP approach (accelerated with the controlled bond expansion) to perform the backward and forward time evolution to the former and the latter, which results in the two states $e^{iHt/2}c_{\mathbb{k}\alpha}^{\dagger}|{\rm GS}\rangle$ and $e^{-iHt/2}c_{\mathbb{k}\beta}^{\dagger}|{\rm GS}\rangle$, respectively. The inner product is calculated after each time step and the sequence of $GS\expectationvalue{c_{\mathbb{k}\alpha}e^{-iHt}c_{\mathbb{k}\beta}^{\dagger}}{\rm GS}$ is collected to be used in the numerical integration latter. Regarding $GS\expectationvalue{c_{\mathbb{k}\alpha}e^{-iHt}c_{\mathbb{k}\beta}^{\dagger}}{\rm GS}$ in the second term of Eq. (S13), the procedure is similar except that the creation and annihilation operators are exchanged.
@@
[…]

### v3

@@ Acknowledgments
Jialin Chen, Qiaoyi Li, and Wei Li are indebted to Bin-Bin Chen and Hongyu Lu for stimulating discussions. The iDMRG calculation is based on TeNPy [61]. The dynamical and finite-temperature calculations are based on TensorKit.jl [87] and FiniteMPS.jl (https://github.com/Qiaoyi-Li/FiniteMPS.jl). This work was supported by the National Natural Science Foundation of China (12222412, 12047503), National Key Projects for Research and Development of China (2024YFA1 409200), Innovation Program for Quantum Science and Technology (2021ZD0301900), Strategic Priority Research Program of Chinese Academy of Sciences (XDB1270100), and the Postdoctoral Fellowship Program of CPSF (GZB202 40772). We thank the HPC-ITP for the technical support and generous allocation of CPU time. Xiaoyu Wang acknowledges financial support from the National High Magnetic Field Laboratory through National Science Foundation (NSF) Grant No. DMR-2128556, and the State of Florida.
Author contributions
Wei Li and Xiaoyu Wang initiated this work. Xiaoyu Wang constructed the lattice model. Wei Li, Jialin Chen, and Qiaoyi Li conducted the model analysis and tensor-network calculations. All authors contributed to the analysis of the results and the preparation of the draft.
@@ I.1 Continuum model of twisted bilayer MoTe2
Monolayer MoTe2 features strong spin-orbit coupling [56], such that the spin ($\uparrow,\downarrow$}) and valley ($\mathbf{K},\mathbf{K}^{\prime}$}) quantum numbers are locked. We hereby use the convention of $\{\mathbf{K},\uparrow\}$ and $\{\mathbf{K}^{\prime},\downarrow\}$ to label the electronic states.
@@
Here $\psi^{\dagger}_{l\uparrow}(\mathbf{r})$ creates an electron in layer $l=b,t$ , with spin $\uparrow$ and in valley $\mathbf{K}$, and at position $\mathbf{r}$. For convenience we drop the valley subscript and label the electrons by their spin quantum number. ${h}_{l}({\mathbf{p}})=-({\mathbf{p}}-\hbar\mathbf{K}_{l})^{2}/2m^{*}$ is kinetic energy, with ${\mathbf{p}}=-i\hbar\nabla_{\mathbf{r}}$ and effective mass $m^{*}$. $\mathbf{K}_{l}$ points sit at the corners of the hexagonal moiré Brillouin zone (see Fig. S1(a) inset). We describe the effect of displacement field $\mathbf{D}$ by adding a potential difference $u_{D}$ between two layers. $\Delta_{b(t)}(\mathbf{r})=2V\sum_{j=1,3,5}\cos(\mathbf{G}_{j}\cdot\mathbf{r}\pm\phi)$ is the intralayer potential for bottom (top) layer, and $\Delta_{T}(\mathbf{r})=w(1+e^{-i\mathbf{G}_{2}\cdot\mathbf{r}}+e^{-i\mathbf{G}_{3}\cdot\mathbf{r}})$ is interlayer tunneling potential. $\mathbf{G}_{i=1,\dots 6}$ are the six smallest reciprocal lattice vectors, with $\mathbf{G}_{1}=\frac{4\pi}{\sqrt{3}L_{M}}\hat{x}$ along the +x+x direction, and $\mathbf{G}_{i+1}$ is the vector obtained by rotating $\mathbf{G}_{i}$ $60^{\circ}$ counterclockwise. Here $L_{M}\approx a_{0}/\theta$ is the moiré unit vector length, with $\theta$ the twist angle and $a_{0}$ the MoTe2 lattice spacing.
$a_{0}$ [Å] $m^{*}$ $m_{e}$] $V$ [meV] $\phi$ $\omega$ [meV]
Ref. [6] 3.47 0.62 8.0 -$89.6^{\circ}$ -8.5
Ref. [54] 3.55 0.62 17.0 $107.7^{\circ}$ -16.0
Ref. [19] 3.52 0.6 20.8 $107.7^{\circ}$ -23.8
Ref. [20] 3.52 0.62 7.5 -$00^{\circ} -11$.3
@@ I.2 Tight binding fit to the continuum model
With the exception of Ref. [20], other DFT calculations predict the upper two moiré bands of a given valley carry Chern numbers +1+1 and 1-1 respectively, including the one we used for this work. This allows us to construct exponentially localized Wannier orbitals without suffering from topological obstruction of a band composite with a finite Chern number [57]. Using the trial state procedure discussed in Ref. [57], we construct exponentially localized Wannier orbitals to describe the Hilbert space spanned by the upper two valence bands per valley, see Fig. S1(b). The two Wannier orbitals per valley reside predominantly on the honeycomb lattice, which is the dual lattice of the moiré triangular lattice made of AA stackings. We hereby label these states as $w_{A}(\mathbf{r})$ and $w_{B}(\mathbf{r})$ respectively, with the subscripts $A,B$ denoting the center of the Wannier orbitals. Separately, these two orbitals are layer polarized, with $w_{A(B)}(\mathbf{r})$ having most orbital weights on the bottom (top) layer.
@@ I.4 Particle-hole transformation
Alternatively, we can perform a particle-hole (PH) transformation $d^{\dagger}_{i\sigma}=c_{i\sigma}$ which inverts the non-interacting band structure (as well as the hopping parameters and chemical potential), while simultaneously we preserve the same functional form for the projected density density interactions but with $n_{i\sigma}=d^{\dagger}_{i\sigma}d_{i\sigma}$. Such a transformation has been used in previous theoretical studies on tMoTe2 (see, e.g., Ref. [20]). Here $|{\Omega}\rangle$ will be a vacuum state for $d$ fermions, and therefore the density-density interaction will not lead to band renormalizations.
Therefore, instead of stating our results for hole fillings $1\leq\nu_{h}<0$, we report results for electron fillings $0<\nu\leq 1$. Accordingly, we present the band structure of the model described by Eq. (S2) in Fig. S3, which includes hopping terms up to the third nearest neighbors following the PH transformation. Without interactions, electrons only fill the lower band [see, Fig. S3(c)] when filling factor $\nu\leq 1$.
It should be noted that the PH transformation employed in our work serves as a strict mathematical tool for reformulating the Hamiltonian, and it is conceptually distinct from any approximate PH symmetry in the flat bands. Since our narrow-band model inherently lacks PH symmetry, it cannot be mapped to the lowest Landau level. This intrinsic asymmetry naturally explains the strong PH asymmetry observed in experiments. Consistently, our DMRG results display the same trend: as shown in the phase diagrams of Figs. 1 and 6 in the main text and Fig. S8, topological phases (CI, FCI, and QAHC) and FMz states are more robust for $\nu>1/2$ than for their PH counterparts at $1-\nu$. For example, the $\epsilon_{r}$ stability ranges at $\nu=2/3$ and $5/9$ are broader than those at $\nu=1/3$ and $4/9$, respectively. Moreover, robust topological phases are observed at $\nu=4/7$, $4/5$, and $3/5$, but not at the corresponding fillings $\nu=3/7$, $1/5$, and $2/5$.
II Details of Methods
II.1 Model simplification.
Since the onsite Coulomb interaction $U$ exceeds the hopping parameters $t_{1},t_{2},t_{3}$ by over an order of magnitude, we can project out double occupancy at low temperatures. The numerical benchmarks between two models are shown in section IV. Then we arrive at a spinful $t$-$V$ model whose Hamiltonian reads
$\begin{split}H=&\sum_{i<j}\big(t_{ij}P_{i}c^{\dagger}_{i\uparrow}c_{j\uparrow}P_{j}+t^{*}_{ij}P_{i}c^{\dagger}_{i\downarrow}c_{j,\downarrow}P_{j}+h.c.\big)-\mu\sum_{i}n_{i}P_{i}+\sum_{i<j}V_{ij}P_{i}n_{i}n_{j}P_{j},\end{split}$ (S12)
In the calculations of fully spin-polarized cases, e.g., FCI and QAHC states, we further simplify the model to a spinless version. This is justified as Eq. (S12) conserves electron numbers for both spin flavors, and the spinless Hamiltonian is given by
$\begin{split}H=&\sum_{i<j}\big(t_{ij}c^{\dagger}_{i}c_{j}+h.c.\big)-\mu\sum_{i}n_{i}+\sum_{i<j}V_{ij}n_{i}n_{j},\end{split}$ (S13)
where $c^{\dagger}_{i}$ creates an electron at site $i$, and $n_{i}=c^{\dagger}_{i}c_{i}$ is the particle number operator. This spinless $t$-$V$ model is employed in the calculation of Hall conductivity, and its validity is confirmed by comparing the results with those from the Hubbard-type model (Eq. (1) in the main text) or the spinful $t$-$V$ model (S12).
The three models and their usage in various simulations are summarized in Table 2.
Model No. $d$ $T=0$ $T>0$ Dynamics
Hubbard-type Eq. (1) 4 $\times$
spinful $t$-$V$ Eq. (S12) 3 $\times$
spinless $t$-$V$ Eq. (S13) 2 $\times$
II.2 Density matrix renormalization group.
II.3 Time-dependent variational principle.
After obtaining the ground state, we employ real-time evolution with the TDVP approach [63, 64] to compute the retarded Green’s function and subsequently derive the single-particle spectral function via Fourier transformation. In the computation of dynamical properties, it is necessary to determine the chemical potential $\mu$ beforehand in order to fix the position of the Fermi level $(\omega=0)$. In the present study, we estimate $\mu$ based on low-temperature calculations (practically at $T\simeq 1.45$K), utilizing the $\mu$-$\nu$ relation (see, e.g., Fig. 6f in the main text). In practice, a Parzen window is applied to suppress non-physical oscillations [65, 66], achieving an energy resolution of $2.67$ meV (with evolution time up to 3 meV-1). We choose time step $\delta t=1/16$ meV-1 and bond dimension $\chi=1000$, which renders accurate results in the TDVP calculations. The data convergence and more technical details are presented in the section V.
II.4 Tangent-space tensor renormalization group.
[…]
III Supplementary ground-state results
III.1 Magnetic gaps for various fillings
@@
III.2 Ground-state results for $\nu=1$
@@
III.3 Ground-state results for fillings $\nu=p/q$ with $q\geq 5$
@@
III.3.1 Charge pumping results
@@ III.3.3 Phase diagram
III.4 Phase diagram for incommensurate filling $\nu\simeq 0.63$
The discovery of QAHCs at incommensurate fillings is a fascinating and significant direction. We have already performed investigations at a filling of $\nu\simeq 0.63$, where such a crystal was recently reported in tMoTe2 [85]. As shown in Fig. S9, using iDMRG simulations at $\nu\simeq 0.63$ on a cylinder with $N_{y}=6$ and varying $\epsilon_{r}\in[13,20]$, we observed a QAHC of $\sigma_{xy}=1$ for $\epsilon_{r}\in[16,20]$ via charge pumping. Further analysis of the charge and hole structure reveals that the hole distribution $h(\bm{r})$ exhibits a pattern consistent with a crystal phase, and the static structure factor $\rho(\bm{q})$ shows peak structures similar to those at $\nu=2/3$. This suggests that the electronic crystal at $\nu\simeq 0.63$ shares a similar origin with that at $\nu=2/3$, as discussed in the main text.
Although our simulations have identified QAHCs at certain incommensurate fillings, establishing their existence over a continuous range of fillings [41] remains challenging. This is primarily due to the significantly larger unit cell sizes required to accurately model such incommensurate systems. Further numerical studies will be necessary to explore this more comprehensively, and we intend to pursue these in future work.
Figure S9: Emergence of QAHC states at incommensurate filling $\nu\simeq 0.63$ on a cylinder of width $N_{y}=6$. a, Hall conductivity $\sigma_{xy}$ as a function of the dielectric constant $\epsilon_{r}$. The presence of QAHC states is highlighted in orange. All these states exhibit ferromagnetic polarization (FMz). b, Charge pumping magnitude $|\Delta Q|$ for various QAHC states, measured after inserting a flux $\Phi_{y}=2\pi$. c, Real-space hole density distribution $h(\bm{r})$, and d, corresponding charge structure factor $\rho(\bm{q})$ for the QAHC state at $\epsilon_{r}=19$. The orange dashed hexagon denotes the first Brillouin zone of the underlying honeycomb lattice. Bond dimension $D=1200$ for all these simulations.
IV Supplementary finite-temperature results
[…]
@@ IV Supplementary finite-temperature results
V computation method and convergence check of the spectral function
@@
[…]
@@
where $\widetilde{W}(\omega)\equiv\int_{-\infty}^{\infty}dt\,e^{i\omega t}W(t)$. Therefore, the effect of a finite $t_{\rm max}$ reduces to the convolution of a window function with the original Green’s function. Directly performing the integration in $0,t_{\rm max}]$ is equivalent to applying a rectangular window function, which leads to strongly non-physical oscillations in frequency domain. In practice, we use the Parzen window function